# CHaRNet: Conditioned Heatmap Regression for Robust Dental Landmark Localization

### v1

@@ Abstract
Identifying anatomical landmarks in 3D dental models is crucial for orthodontic treatment. Manually placing these key points is complex, time-consuming, and requires expert knowledge. While some machine learning methods have been proposed for automatic tooth landmark detection in 3D Intraoral Scans (IOS), research remains limited, with no fully end-to-end approaches that avoid teeth segmentation.
We propose CHaRNet (Conditioned Heatmap Regression Network), the first end-to-end deep learning method for tooth landmark detection in 3D IOS. Unlike traditional two-stage methods that segment teeth before detecting landmarks, CHaRNet directly detects landmarks on the input point cloud. It consists of four key modules: (1) a point cloud encoder, (2) a point cloud decoder with a heatmap regression head, (3) a teeth presence classification head, and (4) the innovative Conditioned Heatmap Regression (CHaR) module. The CHaR module refines landmark regression by leveraging teeth presence classification, enabling dynamic adaptation to cases with missing teeth and improving accuracy in complex dental models.
We evaluate CHaRNet using five point cloud learning algorithms to validate the effectiveness of the CHaR module and test it on a clinical dataset of $1,214$ annotated 3D dental models. Both the dataset and code will be publicly released [^1] Resources will be made public upon acceptance to address the lack of open datasets in orthodontics, promote benchmarking, and inspire new research.
CHaRNet achieves a Mean Euclidean Distance Error (MEDE) of 1.28 mm and a Mean Success Ratio (MSR) of 82.40%, demonstrating robust performance. Notably, it excels in handling irregular dental geometries, such as models with missing teeth. This end-to-end approach streamlines orthodontic workflows, improves 3D IOS analysis precision, and facilitates efficient computer-assisted treatment planning.
Keywords Intraoral Scan (IOS)  $\cdot$ Neural Networks  $\cdot$ Deep Learning  $\cdot$ Landmark Detection  $\cdot$ Orthodontic Treatment Planning
@@ 1 Introduction
Automatic 3D landmark detection plays a crucial role in computer-aided orthodontics, allowing tasks such as orthodontic planning, prosthetic design, and the diagnosis of dental anomalies [1]. Intraoral scans (IOS) are commonly utilized to capture accurate digital surface models of dentition. These scanners represent the 3D surface of teeth, typically as point clouds or meshes. These 3D dental models are highly valuable for simulating procedures such as tooth extraction, movement, deletion, and rearrangement, allowing dentists to predict treatment outcomes more effectively. As a result, digital teeth models have the potential to streamline dentists’ workflows and reduce time spent on labor-intensive tasks.
Existing approaches to perform 3D dental tasks such as teeth segmentation and landmark detection often rely on point cloud-based learning methods [2, 3, 4, 5, 6]. Although effective for general point cloud problems, these methods struggle in dental applications, where incomplete geometries and subtle variations in anatomy, such as missing teeth or the presence of third molars, lack specialized handling [7, 3, 8, 9]. Furthermore, the absolute lack of publicly available dental datasets impedes progress by preventing fair comparisons between methods and limiting the accessibility of this field to researchers outside the specialized domains [8, 9].
In this work, we address these challenges through three key contributions. First, we introduce CHaRNet, an end-to-end deep learning model that directly detects landmarks on the IOS point cloud without relying on previous teeth segmentation, drastically reducing the latency of the overall computation. Second, we propose the Conditioned Heatmap Regression (CHaR) module, a novel enhancement to point cloud learning architectures, which dynamically adapts to missing teeth. Third, we construct and make publicly available a dataset of $1,214$ annotated 3D digital teeth models, which includes diverse complexities, detailed reference landmarks and a taxonomy of dentition types to evaluate models across varying levels of complexity.
Our results highlight the effectiveness of the proposed CHaR module, demonstrating significant performance improvements across multiple point cloud learning models. Notably, the CHaR-augmented PointMLP, referred to as CHaRNet, achieves a Mean Euclidean Distance Error (MEDE) of 0.51 mm on normal dental models and 1.28 mm across all dentition types. It achieves a Mean Success Rate (MSR) of 87.06% on normal dental models and 82.40% across all dentition types. These findings underscore the robustness of CHaRNet, particularly in challenging cases involving several missing teeth.
@@ 2.1 Point cloud learning
Recently, some deep learning (DL) methods operate directly on the 3D point cloud representation of dental models. For example, [33] proposed an end-to-end DL approach based on PointCNN [34] to segment individual teeth and the gingiva from point cloud representation of IOS. [35] made use of FeaStNet [36], a graph convolutional neural network, to also perform teeth segmentation on 3D dental models. The authors in [37] introduced Mask-MCNet to perform instance segmentation of point cloud data from IOS. Mask-MCNet localizes each tooth by predicting its 3D bounding box and simultaneously segments all the points inside each box. TSegNet [38] was proposed as an end-to-end learning-based method for robust tooth segmentation on 3D point cloud dental models. It introduced a distance-aware tooth centroid voting scheme and a confidence-aware cascade segmentation module to handle challenging cases like missing, crowded, or misaligned teeth. DBGANet [39] was proposed as a Dual-Branch Geometric Attention Network for 3D tooth segmentation, utilizing centroid-guided separable attention and Gaussian neighbor attention to capture global geometric structure and refine tooth-gingiva boundaries. DentalMAE [40] extended the self-supervised MeshMAE [41] framework, demonstrating improved generalization and transfer learning for teeth segmentation on 3D intra-oral scans, even with limited training data.
@@
Table 1: Comparison between landmark detection in 3D dental models datasets.
@@
2.2 Landmark detection
Landmark detection is a fundamental task in both computer vision and medical imaging, playing a critical role in tasks such as diagnosis, treatment planning, and surgical simulation. Most classical methods [43, 44, 45, 46] have historically made use of local geometric information to detect landmarks that are highly associated with sharp features. Although these methods are effective within their well-defined domains, they are limited in scope and are unable to identify landmarks outside their domain. With the rise of DL, two approaches dominate the data-driven landmark localization problem. The most intuitive approximation is the regression-based approach, which directly regresses the coordinates from the input (images, volumes, point clouds, etc.) [47, 48, 49, 50]. However, this direct mapping is extremely challenging given the unbounded nature of the coordinates [51]. Instead, heatmap-based detection introduced in [52] has become a popular and preferred approach for landmark detection due to its ability to encode location information as probabilities over the inputs, simplifying the task by reducing the space of possibilities to the number of elements of the input space (pixels, voxels, vertices, etc.) [53, 54, 55].
In recent years, landmark localization task has become popular in medical image analysis [56, 57, 58, 59]. However, only a handful of works have been proposed for landmark detection in 3D dental models. [1] proposed a set of analytical approaches to identify dental-specific features (e.g., cusps, marginal ridges, and grooves) on digital dental meshes. At the individual tooth level, [3] designed a novel neural network architecture for the joint tasks of predicting landmarks and axes. [42] presented a method for landmark detection in 3D dental meshes that leverages a multi-view approach, transforming the task into the 2D domain. In this way, the network detects landmarks through heatmap regression across multiple viewpoints. ALR [60] first determines the orientation of the scan, then uses local maxima along the vertical axis as an initial approximation for landmarks. This is followed by analyzing surface gradient and curvature information to identify the shape and boundaries of each tooth. c-SCN [7] was proposed as an end-to-end method for tooth segmentation and landmark localization but only on teeth crowns. TS-MDL [8] proposed a two-stage framework to subsequently perform teeth segmentation and landmark detection on IOS. By harnessing the segmentation outputs, authors used a variant of PointNet to detect landmarks for each individual tooth. While effective, this approach comes at a high computational cost derived from separating the process into two stages. Besides, the used dataset includes mainly regular dental models without considering complex dental models with missing teeth and one or two third molars. ALIIOS algorithm [9] combines image processing, segmentation, and DL to identify dental landmarks on IOS by synthesizing 3D patches with the output of a 2D U-Net. However, most of these methods rely on a previous segmentation of teeth, which adds a considerable latency to the overall system.
3 Data construction
Existing datasets for landmark detection in 3D intraoral scans are very small, private and do not include frequent anomalies of the dentition. A comparison between these datasets is provided in Table 1. In this work, we construct a comprehensive dataset of $1,214$ digital 3D dental models derived from post-processed raw scans. Each model is annotated with detailed reference landmarks and categorized into a taxonomy of dentition types to facilitate the evaluation of automatic 3D landmark detection systems under varying levels of difficulty. Our dataset does not only provide a robust foundation for training and evaluating point cloud learning methods but also addresses a significant gap in the orthodontic field, where publicly available annotated datasets are scarce.
In this regard, we make our dataset publicly available so that the community can explore new research ideas on it. This is also an attempt to encourage the orthodontic community to publish their datasets and hence allow for a fair comparison between different AI methods, attracting more practitioners to the field of computer-assisted orthodontics, and encouraging the development of new domain-specific technologies.
@@ 3 Data construction
3.1 Reference landmarks
Each dentition in our dataset contains 5 landmarks per tooth, making up to 80 landmarks as showed in Figure 1, depending on the number of teeth present in each dentition. These 5 landmarks represent key points for subsequent tasks, such as treatment planning, are described in Table 2.
Table 2: Description of dental landmarks
@@ 3.3 Teeth distribution
Figure 4: Original (left column) and downsampled (right column) point clouds.
@@ 3.4 Data preprocessing
Many dental models in the dataset have missing teeth, and since we use a heatmap regression approach for landmark detection, it is necessary to account for landmarks associated with these missing teeth. Allocating these landmarks to any point within the original downsampled point cloud would be illogical, as the corresponding teeth are absent. To address this issue, we introduce an additional point located outside the downsampled point cloud, called the "null point" (see Figure 5). This null point serves as a placeholder for the landmarks of missing teeth, ensuring that the model can handle incomplete dental models effectively. By adding the null point, each point cloud will contain $10001$ vertices in total.
@@ 3.5 Ground-truth generation
$h_{ki}(x_{i})=\exp\left(-\frac{d(x_{i},l_{k})^{2}}{2\sigma^{2}}\right)$ (1)
@@ 4.1 Problem statement
Given an input point cloud $P=\{x_{i}\in\mathbb{R}^{3}\mid i=1,...,N\}$ (where $N=10001$, including the "null" point), our task is to predict the positions $L=\{l_{k}\in\mathbb{R}^{3}\mid k=1,...,80\}$ of 80 dental landmarks. Our method employs a heatmap-based approach to encode the likelihood of each point $x_{i}$ being a particular landmark. For each landmark $k$, the heatmap $\hat{H}_{k}=\{\hat{h}_{ki}(x_{i})\mid i=1,...,N\}$ represents these likelihoods. Then, the predicted landmark positions are defined as:
$\hat{L}=\{\hat{l}_{k}=x_{i}\text{ where }\hat{h}_{ki}=\max(\hat{H}_{k})\mid k=1,...,80\}.$ (2)
@@
For the heatmap regression task, we minimize the mean squared error (MSE) loss between the ground truth heatmaps $H_{k}=\{h_{ki}(x_{i})\mid i=1,...,N\}$ and the predicted heatmaps $\hat{H}_{k}$. Specifically, for each landmark $k$, the MSE loss is computed as:
$\mathcal{L}_{\text{MSE}}=\frac{1}{K}\sum_{k=1}^{K}\frac{1}{N}\sum_{i=1}^{N}\left(h_{ki}(x_{i})-\hat{h}_{ki}(x_{i})\right)^{2},$ (3)
@@ 4.2 Network
The point cloud encoder is in charge of extracting a reduced internal representation of the whole input point cloud. Point cloud learning methods usually perform this dimensionality reduction and feature selection hierarchically through set abstraction operations [61, 62]. This learned high-level features will subsequently be fed to the decoder plus heatmap regression head, and the teeth presence classification.
@@
This is achieved by generating a set of heatmaps, where each heatmap corresponds to a landmark and encodes the likelihood distribution over all points in the point cloud. Formally, for each landmark $k$ we obtain the initial heatmap $\hat{H}_{k}^{\prime}=\{\hat{h}_{ki}^{\prime}(x_{i})\mid i=1,...,N\}$, which will be used to compute the final heatmap $\hat{H}_{k}$ as explained below.
@@
$\hat{H}_{t,g}=\{h_{(t,g)i}^{\prime}(x_{i})\cdot p_{t}\mid i=1,...,N-1\}\cup\{h_{(t,g)N}^{\prime}(x_{N})\cdot(1-p_{t})\}$ (6)
In this formulation, the weights $p_{t}$ and $(1-p_{t})$ act as a switch: when $p_{t}$ is high (tooth $t$ is present), the likelihoods of points in the original point cloud dominate, encouraging the landmarks associated with tooth $t$ to remain in the point cloud. In contrast, when $p_{t}$ is low (tooth $t$ is absent), the probability of a null point dominates, shifting the associated landmarks to the null point. Figure 7 provides a schematic representation of our CHaR module. For clarity, we denote the set of likelihood values for the original points (excluding the null point) as $\mathbf{h}_{(t,g)}=\{h_{(t,g)i}^{\prime}(x_{i})\mid i=1,...,N-1\}$, and the likelihood value for the null point as $n_{(t,g)}=\{h_{(t,g)N}^{\prime}(x_{N})\}$.

### v4

@@ Abstract
Identifying anatomical landmarks in 3D dental models is vital for orthodontic treatment, yet manual placement is complex and time-consuming. Although some machine learning approaches have been proposed for automatic tooth landmark detection in 3D Intraoral Scans (IOS), none provide a fully end-to-end solution that bypasses teeth segmentation, limiting practical applicability.
We introduce CHaRNet (Conditioned Heatmap Regression Network), the first fully end-to-end deep learning framework for tooth landmark detection in 3D IOS. Unlike traditional two-stage workflows that segment teeth before detecting landmarks, CHaRNet directly operates on the input point cloud, thus reducing complexity and computational overhead. Our method integrates four modules: (1) a point cloud encoder, (2) a point cloud decoder with a heatmap regression head, (3) a teeth presence classification head, and (4) the novel Conditioned Heatmap Regression (CHaR) module. By leveraging teeth presence classification, the CHaR module dynamically adapts to missing teeth and enhances detection accuracy in complex dental models.
We evaluate CHaRNet using five point cloud learning algorithms on a clinical dataset of 1,214 annotated 3D models. Both the dataset and code will be publicly released [^1] Resources will be made public upon acceptance to address the lack of open datasets in orthodontics and inspire further research. CHaRNet achieves a Mean Euclidean Distance Error (MEDE) of 0.51 mm on typical dental models and 1.28 mm across all dentition types, with corresponding Mean Success Rates (MSR) of 87.06% and 82.40%, respectively. Notably, it exhibits robust performance on irregular geometries, including models with missing teeth. This end-to-end approach streamlines orthodontic workflows, enhances 3D IOS analysis precision, and supports efficient computer-assisted treatment planning.
Keywords Intraoral Scan  $\cdot$ Neural Networks  $\cdot$ Deep Learning  $\cdot$ Landmark Detection  $\cdot$ Orthodontic Treatment Planning
@@ 1 Introduction
Automatic 3D anatomical keypoints or landmarks detection plays a crucial role in computer-aided orthodontics, allowing tasks such as orthodontic planning, prosthetic design, and the diagnosis of dental anomalies [1]. Intraoral scans (IOS) are commonly utilized to capture accurate digital surface models of dentition. These scanners represent the 3D surface of teeth, typically as point clouds or meshes. These 3D dental models are highly valuable for simulating procedures such as tooth extraction, movement, deletion, and rearrangement, allowing dentists to predict treatment outcomes more effectively. As a result, digital teeth models have the potential to streamline dentists’ workflows and reduce the time spent on labor-intensive tasks.
Existing approaches to perform 3D dental tasks such as teeth segmentation and landmark detection often rely on point cloud-based learning methods [2, 3, 4, 5, 6]. Although effective for general point cloud problems, these methods struggle in dental applications, where incomplete geometries and subtle variations in anatomy, such as missing teeth or the presence of third molars, lack specialized handling [7, 3, 8, 9]. Furthermore, the absolute lack of publicly available dental datasets impedes progress by preventing fair comparisons between methods and limiting accessibility to this field to researchers outside the specialized domains [8, 9].
In this work, we address these challenges through three key contributions:
  1. 1.
    We introduce CHaRNet, an end-to-end deep learning (DL) model that directly detects landmarks on the IOS point cloud without relying on previous teeth segmentation, drastically reducing the latency of the overall computation.
  2. 2.
    We propose the Conditioned Heatmap Regression (CHaR) module, a novel enhancement to point cloud learning architectures, which dynamically adapts to missing teeth.
  3. 3.
    We construct and make publicly available a dataset of $1,214$ annotated 3D digital teeth models, which includes diverse complexities, detailed reference landmarks and a taxonomy of dentition types to evaluate models across varying levels of complexity.
Our results highlight the effectiveness of the proposed CHaR module, demonstrating significant performance improvements across multiple point cloud learning models. Notably, the CHaR-augmented PointMLP, referred to as CHaRNet, achieves a Mean Euclidean Distance Error (MEDE) of 0.51 mm on typical dental models, defined as those with a complete set of teeth excluding third molars, and 1.28 mm across all dentition types. It achieves a Mean Success Rate (MSR) of 87.06% on typical dental models and 82.40% across all dentition types. These findings underscore the robustness of CHaRNet, particularly in challenging cases involving several missing teeth.
@@ 2.1 Point cloud learning
Recently, some DL methods operate directly on the 3D point cloud representation of dental models. For example, [33] proposed an end-to-end DL approach based on PointCNN [34] to segment individual teeth and gingiva from point cloud representation of IOS. [35] made use of FeaStNet [36], a graph convolutional neural network, to also perform teeth segmentation on 3D dental models. The authors in [37] introduced Mask-MCNet to perform instance segmentation of point cloud data from IOS. Mask-MCNet localizes each tooth by predicting its 3D bounding box and simultaneously segments all the points inside each box. TSegNet [38] was proposed as an end-to-end learning-based method for robust tooth segmentation on 3D point cloud dental models. It introduced a distance-aware tooth centroid voting scheme and a confidence-aware cascade segmentation module to handle challenging cases like missing, crowded, or misaligned teeth. DBGANet [39] was proposed as a Dual-Branch Geometric Attention Network for 3D tooth segmentation, utilizing centroid-guided separable attention and Gaussian neighbor attention to capture global geometric structure and refine tooth-gingiva boundaries. DentalMAE [40] extended the self-supervised MeshMAE [41] framework, demonstrating improved generalization and transfer learning for teeth segmentation on 3D intra-oral scans, even with limited training data.
2.2 Landmark detection
Landmark detection is a fundamental task in both computer vision and medical imaging, playing a critical role in tasks such as diagnosis, treatment planning, and surgical simulation. Most classical methods [42, 43, 44, 45] have historically made use of local geometric information to detect landmarks that are highly associated with sharp features. Although these methods are effective within their well-defined domains, they are limited in scope and are unable to identify landmarks outside their domain. With the rise of DL, two approaches dominate the data-driven landmark localization problem. The most intuitive approximation is the regression-based approach, which directly regresses coordinates from the input (images, volumes, point clouds, etc.) [46, 47, 48, 49]. However, this direct mapping is extremely challenging given the unbounded nature of the coordinates [50]. Instead, heatmap-based detection, introduced in [51], has become a popular and preferred approach for landmark detection due to its ability to encode location information as probabilities over the inputs. This approach simplifies the task by reducing the space of possibilities to the number of elements in the input space (e.g., pixels, voxels, vertices) [52, 53, 54].
In recent years, landmark localization task has become popular in medical image analysis [55, 56, 57, 58]. However, only a handful of works have been proposed for landmark detection in 3D dental models. [1] proposed a set of analytical approaches to identify dental-specific features (e.g., cusps, marginal ridges, and grooves) on digital dental meshes. At individual tooth level, [3] designed a novel neural network architecture for the joint tasks of predicting landmarks and axes. [59] presented a method for landmark detection in 3D dental meshes that leverages a multi-view approach, transforming the task into the 2D domain. In this way, the network detects landmarks through heatmap regression across multiple viewpoints. ALR [60] first determines the orientation of the scan, then uses local maxima along the vertical axis as an initial approximation for landmarks. This is followed by analyzing surface gradient and curvature information to identify the shape and boundaries of each tooth. c-SCN [7] was proposed as an end-to-end method for tooth segmentation and landmark localization but only on teeth crowns. TS-MDL [8] proposed a two-stage framework to subsequently perform teeth segmentation and landmark detection on IOS. By harnessing the segmentation outputs, authors used a variant of PointNet to detect landmarks for each individual tooth. While effective, this approach comes at a high computational cost derived from separating the process into two stages. Besides, the used dataset includes mainly regular dental models without considering complex dental models with missing teeth and one or two third molars. ALIIOS algorithm [9] combines image processing, segmentation, and DL to identify dental landmarks on IOS by synthesizing 3D patches with the output of a 2D U-Net. However, most of these methods rely on a previous segmentation of teeth, which adds considerable latency to the overall system.
@@ 2.2 Landmark detection
Table 1: Comparison between datasets for landmark detection in 3D dental models.
@@
Table 2: Description of dental landmarks.
@@
3 Data construction
Existing datasets for landmark detection in 3D intraoral scans are very small, private and do not include frequent anomalies of the dentition. A comparison between these datasets is provided in Table 1. In this work, we construct a comprehensive dataset of 1,214 digital 3D dental models derived from post-processed raw scans. Each model is annotated with detailed reference landmarks and categorized into a taxonomy of dentition types to facilitate the evaluation of automatic 3D landmark detection systems under varying levels of difficulty. Our dataset does not only provide a robust foundation for training and evaluating point cloud learning methods but also addresses a significant gap in the orthodontic field, where publicly available annotated datasets are scarce. In this regard, we make our dataset publicly available so that the community can explore new research ideas on it. This is also an attempt to encourage the orthodontic community to publish their datasets and hence allow for a fair comparison between different AI methods, attracting more practitioners to the field of computer-assisted orthodontics, and encouraging the development of new domain-specific technologies.
To enable accurate and consistent annotations, each dentition in our dataset contains five landmarks per tooth, making up to 80 landmarks in total (depending on the number of teeth present in each dentition). These landmarks represent key points for subsequent tasks, such as treatment planning, and are illustrated in Figure 1 and detailed in Table 2. Specifically, for each tooth, the five landmarks identify important anatomical structures that help guide orthodontic measurements and interventions. This meticulous labeling ensures that every model in the dataset captures essential dental characteristics, ranging from normal anatomical variations to more complex anomalies, thereby broadening the scope and applicability of the dataset for diverse research and clinical needs.
@@ 3.2 Teeth distribution
Figure 4: Original (top row) and downsampled (bottom row) point clouds.
@@ 3.3 Data preprocessing
Many dental models in the dataset have missing teeth, and since we use a heatmap regression approach for landmark detection, it is necessary to account for landmarks associated with these missing teeth. Allocating these landmarks to any point within the original downsampled point cloud would be illogical, as the corresponding teeth are absent. To address this issue, we introduce an additional point located outside the downsampled point cloud, called the "null point" (see Figure 5). This point serves as a placeholder for the landmarks of missing teeth, ensuring that the model can handle incomplete dental models effectively. The location of the null point is determined in a data-dependent manner and is computed as follows: Let $m_{b}$ represent the largest extent of the bounding box enclosing the point cloud across the three axes, and let $c$ denote the centroid of the point cloud. The null point $n$ is then placed at $n=c+\frac{m_{b}}{2}\cdot(0,1,0)$, ensuring it is positioned outside the point cloud along the positive $y$-axis. By adding this point, each point cloud will contain $10001$ vertices in total.
@@ 3.4 Ground-truth generation
$h_{ki}=\exp\left(-\frac{d(x_{i},l_{k})^{2}}{2\sigma^{2}}\right)$ (1)
@@ 4.1 Problem statement
Given an input point cloud $P=\{x_{i}\in\mathbb{R}^{3}\mid i=1,...,N\}$ (where $N=10001$, including the "null" point), our task is to predict the positions $L=\{l_{k}\in\mathbb{R}^{3}\mid k=1,...,80\}$ of 80 dental landmarks. Our method employs a heatmap-based approach to encode the likelihood of each point $x_{i}$ being a particular landmark. For each landmark $k$, the heatmap $\hat{H}_{k}=\{\hat{h}_{ki}\mid i=1,...,N\}$ represents these likelihoods. Then, the predicted landmark positions are defined as:
$\hat{L}=\{\hat{l}_{k}=\underset{x_{i}\in P}{\operatorname{argmax}}(\hat{H}_{k})\mid k=1,...,80\}.$ (2)
@@
For the heatmap regression task, we minimize the mean squared error (MSE) loss between the ground truth heatmaps $H_{k}=\{h_{ki}\mid i=1,...,N\}$ and the predicted heatmaps $\hat{H}_{k}$. Specifically, for each landmark $k$, the MSE loss is computed as:
$\mathcal{L}_{\text{MSE}}=\frac{1}{K}\sum_{k=1}^{K}\frac{1}{N}\sum_{i=1}^{N}\left(h_{ki}-\hat{h}_{ki}\right)^{2},$ (3)
@@ 4.2 Network
The point cloud encoder is in charge of extracting a reduced internal representation of the whole input point cloud. Point cloud learning methods usually perform this dimensionality reduction and feature selection hierarchically through set abstraction operations [PointNet2, PointNext]. This learned high-level features will subsequently be fed to the decoder plus heatmap regression head, and the teeth presence classification.
@@
This is achieved by generating a set of heatmaps, where each heatmap corresponds to a landmark and encodes the likelihood distribution over all points in the point cloud. Formally, for each landmark $k$ we obtain the initial heatmap $\hat{H}_{k}^{\prime}=\{\hat{h}_{ki}^{\prime}\mid i=1,...,N\}$, which will be used to compute the final heatmap $\hat{H}_{k}$ as explained below.
@@
$\hat{H}_{t,g}=\{h_{(t,g)i}^{\prime}\cdot p_{t}\mid i=1,...,N-1\}\cup\{h_{(t,g)N}^{\prime}\cdot(1-p_{t})\}$ (6)
In this formulation, the weights $p_{t}$ and $(1-p_{t})$ act as a switch: when $p_{t}$ is high (tooth $t$ is present), the likelihoods of points in the original point cloud dominate, encouraging the landmarks associated with tooth $t$ to remain in the point cloud. In contrast, when $p_{t}$ is low (tooth $t$ is absent), the probability of a null point dominates, shifting the associated landmarks to the null point. Figure 7 provides a schematic representation of our CHaR module. For clarity, we denote the set of likelihood values for the original points (excluding the null point) as $\mathbf{h}_{(t,g)}=\{h_{(t,g)i}^{\prime}\mid i=1,...,N-1\}$, and the likelihood value for the null point as $n_{(t,g)}=h_{(t,g)N}^{\prime}$.